OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x - 2y = 8. Answer x intercept = (-8, 0); y intercept = (0, -4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (8, 0); y intercept = (0, 4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (aravindg):

divide whole equation by 8

OpenStudy (aravindg):

\[\huge \frac{x}{8}+\frac{y}{-4}=1\] which means x intercept=8 and y intercept=-4

OpenStudy (aravindg):

that will be option 2

OpenStudy (aravindg):

x intercept = (8, 0); y intercept = (0, -4)

OpenStudy (aravindg):

any doubts?

OpenStudy (aravindg):

:)